A car moves from rest on a journey with an acceleration of 2m/s^2 for 2sec,and continues with a uniform velocity for further 8sec and then finally came to retardation and stop in 8sec.calculate the maximum velocity attained during this journey.(a) draw a velocity time graph of the journey (b) calculate the distance travelled during the first 12sec (c)calculate the distance travel (d)calculate the average retardation?
1
Expert's answer
2020-02-24T10:57:52-0500
"v=4\frac{m}{s}"

a)



b)


"v(12)=4-0.5(2)=3\frac{m}{s}"


"d(12)=0.5(4)(2)+(4)(8)+0.5(4+3)(2)=43\ m"

c)


"d(18)=0.5(4)(2)+(4)(8)+0.5(4)(8)=52\ m"

d)


"a=\frac{4}{8}=0.5\frac{m}{s^2}"